Consider a hypothetical population with two genes. Each of these two genes have two alleles (A & a, B & b). These genes are not linked, and these genes are true dominant/recessive (i.e. not co-dominant or incomplete dominance... "A" dominant to "a" and "B" dominant to "b"). Suppose you perform a monohybrid cross of an AABB homozygote with an aabb homozygote through to the F2 generation. In 2019, the genotype frequency of individuals with "aa" (regardless of their B/b genes) was at 0.09 and the population reproduces once per year. If in 2020, there are 40 heterozygotes (Aa genotypes) out of a total population of 80 individuals, what can we say?

Group of answer choices

A) The population didn't evolve

B) The population evolved via gene flow

C) The population evolved via genetic drift

D) The population evolved, but we can't say why
Answer #1

D. Population evolved, but we can't say why.

This is the right option.

In 2019, aa frequency (q^2) is 0.09.

q = square root of 0.09 = 0.3

As per Hardy -Weinberg equation-- p = q = 1

So, p = 1- q = 1- 0.3 = 0.7

Heterozygote frequency ( 2pq ) in 2019 is --- 0. 3 x 0.7 x 2 = 0.42

In 2020 heterozygote frequency is = heterozygotes number / total offspring

2 pq = 40 / 80 = 0. 5

In 2019 the heterozygote frequency is 0.42 and in 2020, heterozygote frequency is 0.5.

Change in allele frequency is evolution. So in this population , evolution is occurring. But why it is occurring is not known.